(3-2i)(2+4i)
please show work

Answer by Alan3354(69443)

(3-2i)(2+4i)
= 3(2+4i) - 2i(2+4i)
= 6+12i - 4i - 8i*i
(i*i = -1), so -8i*i = +8
= 6+12i - 4i + 8
= 14 + 8i